OpenStudy (anonymous):

A polygon has an area of 100 square inches and one of its sides is 6 inches long. If a second similar polygon has an area of 36 square inches, what is the length of the corresponding side in the second polygon?

OpenStudy (anonymous):

the ratio of corresponding side = sqrt 100 / sqrt 36 = 10 / 6 = 5 / 3

OpenStudy (anonymous):

can you continue from this?

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

@joeywhite, can you explain how you arrived at that ratio?

OpenStudy (anonymous):

yes the ratio of the area of 2 similar polygons are the ratio of the squares of corresponding sides so ratio of corresponding sides = ratio of square roots of areas

OpenStudy (anonymous):